Calculate the hydronium ion, H3O+ , and hydroxide ion, OH− , concentrations for a 0.0111 M HCl solution.

[H3O+]=

[OH−]= 




1
Expert's answer
2022-04-22T04:36:03-0400

Solution:

HCl is a strong acid that completely dissociates into hydronium (H3O+) and chloride (Cl) ions.

The dissociation equation for HCl in aqueous solution is:

HCl(aq) + H2O(l) → H3O+(aq) + Cl(aq)

From the reaction, one hydronium ion and one chloride ions are formed.

So, the concentration of H3O+ is same as the concentration of HCl.

Therefore,

[H3O+] = 0.0111 M


Kw = [H3O+] × [OH] = 1.0×10−14

Therefore,

[OH] = Kw / [H3O+] = (1.0×10−14) / (0.0111) = 9.01×10−13

[OH] = 9.01×10−13 M


Answer:

[H3O+] = 0.0111 M

[OH] = 9.01×10−13 M